Today's shift was devoted to two activities:
- study of non-linear effects, by increasing particular injected lines and observing effects on hrec;
- update of the main ISC processes (LSC_Acl, LSC_Acl_moni, ASC_Acl and ASC_pre) with the introduction of a new DARM line at low frequency (33.7 Hz) to be used for a tentatively better signal for the OS/SRCL_SET loop; additionally, we wanted to get rid of the DARM_DCP_LF line at 176.3 Hz and replace it with the 74.4 Hz which is always on anyway; additionally, we wanted a new set of SR AA signals which use the same technique of the current one (DARM double demodulation and phi extraction), but demodulated at 74.4 Hz instead of 491.3 Hz.
Study of non-linear effects:
We started with the ITF in LOW_NOISE_3; we opened the SR loop and misaligned SR (+0.7 urad TY) to null the CMRF LF and maximize the BNS Range; then we injected back the SRCL line (off at that point), increasing its amplitude over several steps:
8:49 UTC 120 sec SRCL line ampl 0
8:52:0UTC 120 sec SRCL line ampl 0.001
8:56 UTC 120 sec SRCL line ampl 0.002
9:00 UTC 120 sec SRCL line ampl 0.004
9:03 UTC 120 sec SRCL line ampl 0.008
9:07 UTC 120 sec SRCL line ampl 0.016
9:10 UTC 120 sec SRCL line ampl 0.032
9:13 UTC 120 sec SRCL line ampl 0.05
Since the line is visible in hrec we could see already it induces other peaks (see Figure); more analysis will follow.
Then we closed again SR to restore its initial aligned position, then we misaligned again with the same purpose (although this time the best Range was achieved at +1.1 urad TY), then we started with the DIFFp TX line; given its low frequency (4.8 Hz) it is not visible directly in hrec and it will need offline analysis:
10:07:00 UTC SR AA open;
10:10 UTC to 10:32 UTC: SR misalignment (+1.1 urad TY) and ITF stabilization;
10:32 to 10:37 UTC: progressively increased DIFFp TX line amplitude;
10:37:00 UTC : 180 s with DIFFp TX line at 1e-3 amplitude, which is 10x the amplitude we have in LOW_NOISE_2 (in LOW_NOISE_3 is turned off completely).
Unfortunately we unlocked before moving to TY, so we moved to the second topic.
New DARM line at 33.7 Hz and restart of ISC processes:
We then moved to complete the work already started a while back, adding/changing a few things with respect to the code already deployed offline.
An initial disclaimer about nomenclature, which I think has become a bit obscure over time to some:
- with "DARM line" we intend the main one, used for a number of things, which is injected at 74.4 Hz;
- back in the day we added two lines to be used for the DCP computation, although this changed over time; these are called "DARM LF line" (176.3 Hz) and "DARM HF line" (491.3 Hz) as they were intended as lower than the DCP and higher than the DCP respectively; the former is now unused, not deleted, and kept to zero; the latter is used for many important things, and has been left unchanged by the work described here and in the referenced entry;
- we introduce now a new "DARM LLF line", injected at 33.7 Hz; its purpose is to hopefully deliver a better signal for the OS, as it is closer to where the OS actually is.
In order not to have too many lines, as we introduced a new one we wanted to get rid of the DARM LF line, as it is used only for the DCP monitor, evaluated as the ratio of the mag of the demodulation of DARM at the HF and the LF line frequencies; now, instead, we replace the 176.3 Hz line with the 74.4 Hz one, which is always on as it is used also for the DARM UGF and other things.
So, this is the recap of the work done:
- Optical Spring / SRCL_SET: with the new LLF line, we duplicated the whole OS code (_LLF_ version of the same channels), so now we have the same monitor as before, plus the new one, which is online but not calibrated just yet; for this reason, we changed a bit the code for the SRCL_SET servo, and now SRCL_SET_INPUT has become a weighted sum (ACL_SUM_CH object) of the two signals provided by the two machineries; right now, the usual signal has weight 1 and the new one weight 0; both of them have their calibration upstream with respect to such weight, and those would be the ones to be measured and set; a new set of demodulation filters at the LLF frequency have been added to LSC_Acl as well;
- Double-Cavity Pole monitor: as explained above, now the LSC_DCP_moni_mad_cal channels is computed as the ratio of the magnitudes of the DARM demodulations at 491.Hz and 74.4 Hz, where the latter replaced the demodulation at 176.3 Hz; the calibration in CARM_NULL_1F (which was reasonably correct in Hz) has not changed considerably, at least at order zero; instead, the calibration in LOW_NOISE_2 (which was already incorrect in Hz, as it peaked at ~ 3000) has strongly changed, and now a high value of the DCP (cross-checked with the CAL estimator) corresponds to ~ 98 in those arbitrary units;
- SR alignment signals: we added a new set of signals for the SR alignment; these use the same technique as the current ones, which means a double demodulation of DARM at a DARM line (phi extraction) and then at a SR TX/TY line; the DARM demodulation is done at 74.4 Hz (instead of 491.3 Hz) and as such they should be more sensitive to the DARM optical gain rather then the DCP; these signals are named SR_T*_DCP_phi_B1_I and are generated in ASC_pre as the other ones; for internal consistency, though, once they are acquired in ASC_Acl they get the name the current one have, and the current ones have an additional _HF_ segment in their name to reflect the HF demodulation instead of the 74.4 Hz one. This should be transparent, though, as those duplicated were and are not saved in the DAQ, and the aforementioned channels from ASC_pre should be used instead, as usual. The new channels are online but they are still missing demodulation phase and calibration.
After the restart of the processes, we verified that the OS demodulation phase was still good, we retuned the demodulation phases for the SR alignment signals (both the B1p- and the B1-based ones), and we relocked up to LOW_NOISE_2.
The Automation has been already update to reflect all the changes reported above; in particular:
- the new DARM LLF line is reset in DOWN, and it is turned on during lock acquisition; its amplitude is almost double the one of the 74.4 Hz one in CARM_NULL_1F, and equal to that in later states; differently from the 74.4 Hz line, it is not turned on before CARM_NULL_1F;
- duplicated commands for the new OS monitor have been changed, numerical values still need to be updated, following calibration;
- the sensing matrix for the SR alignment already reflects the changes in the SR signals, and a new section will need to be uncommented (commenting the current one) in case of use; numerical parameters are there but will need measurements.